Two race cars, A and B, have numbers that are consecutive odd integers whose sum is 88. If A is smaller, find the number on each car.
Answer by KnightOwlTutor(293)

X=A 1rst odd number B=X+2=2nd consecutive odd number
X+X+2=88
2X+2=88
subtract 2 from both sides
2X=86
X=43
X+2=45
Check by plugging in values
43+45=88